I included all of the information the book provided. I found the textbook solution but I need help determining which numbers get plugged in to populate the answer. If I can understand how the return on equity, return on assets, ect is computed on one I can solve the remaining years/companies. I keep getting answers one over.

Question a- Compete key liquidity, solvency, and return on investment ratios for 1998 (current ratio, total debt to equity, long-term debt to equity, times interests earned, return on assets, return on equity).

I cannot determine which numbers are used to complete the equation.on the balance sheet, mos erm, as operating leases most leases. The md rowing in allere volatile carmings Airline companies lUAB DELTA 1997 AMR 1998 1998 1997 1998 1997 $ 1 $ $ 5,077 Dies and equity $ 176 5,492 $ 62 4021 $ 4.514 154 5.485 135 5.437 e(concluded) DELTA AMR VAL Operating Capital Operating Capital Opera Operating Much of pers Morris $101 $ 288 $ 1,011 985 $ 14
Answer #1
AMR
Ratio Formulae 1998
Current ratio = Current assets / Current liabilities =4875/(154+5485) = 0.86
Total debt to equity = (Short term debt + Long term debt) / Shareholder equity =(154+2436)/(3257+4729-1288) = 0.39
Long-term debt to equity = Long term debt / Shareholder equity =(2436)/(3257+4729+(-1288)) = 0.36
Times interests earned = Operating income / Interest expense =2338/372 = 6.28
Return on assets = Net income / Total assets =1306/22303 = 5.86%
Return on equity = Net income / Shareholders equity =1306/(3257+4729+(-1288)) = 19.50%
Note : Shareholders equity = Contributed capital + Retained earnings + Treasury stock